Please help me solve this equation by using the distributive property to rewrite without parentheses. -y (-y+y)
Answer by stanbon(75887)

using the distributive property to rewrite without parentheses.
-y (-y+y) = -y(-y)+(-y(y)) = y^2 - y^2 = 0
=======================
Cheers,
Stan H.
------------